### v1

@@ Appendix 0.B Concrete Usecase: Uniswap
Uniswap has three versions: Uniswap V1, released in November 2018, consisted of the baseline protocol that implemented ERC20 token swaps with Ethereum and all of the liquidity management methods (mint, burn, collect). Uniswap V2, released in August 2020, introduced ERC20 to ERC20 swaps, liquidity provision incentives, and oracles. And Uniswap V3 released in May 2021, introduced concentrated liquidity, a nonfungible representation of liquidity positions, and further improvements to the oracle systems. Uniswap adopts the constant product formula for computing the trading price described in Section 2. As discussed in Section 1, Uniswap is among the most popular AMMs in practice and commands a large market share in the AMM industry. In this section, we provide an overview of the set of contracts that implement the Uniswap functionality on Ethereum, and the execution trace of the supported transactions.
@@ 0.B.1 Uniswap Supporting Contracts
Pool factory and deployer. The $\mathsf{PoolFactory}$ and $\mathsf{PoolDeployer}$ contracts are responsible for setting up new token pools. The pool deployer contract provides the interface, and the pool factory contract performs the heavy lifting of the deployment process, i.e., creating the actual pool. Once a pool is created, clients and LPs can start interacting with the pool.
@@ 0.B.2 Transaction Execution Trace
Swap. Regardless of the type of swap (exact in/out) being executed, clients must first deposit their input tokens in the $\mathsf{SwapRouter}$ contract, and approve it to spend their tokens. The client then calls the relevant function of the $\mathsf{SwapRouter}$ contract ($\mathsf{ExactInput}$ or $\mathsf{ExactOutput}$) to submit a swap transaction. If the user is performing an $\mathsf{ExactOutput}$ swap, they should implement an additional set of conditional transfers to occur after the call to $\mathsf{ExactOutput}$ to retrieve unspent input tokens.
@@
Flash. To execute a flash transaction, the client must begin by writing and deploying a smart contract which overwrites the $\mathsf{Flash}_{callback}$ method of the liquidity pool. The $\mathsf{Flash}_{callback}$ function is responsible for paying back the loan. As such, should a client want to perform arbitrage with the loan, they begin by overwriting $\mathsf{Flash}_{callback}$, simply adding in solidity code to execute their arbitrage opportunity. They can then call the flash function of the liquidity pool from which they would like to execute a flash transaction. $\mathsf{Flash}$ itself simply transfers the requested loan of tokens to the client, where they are used for the arbitrage opportunity, before being re-transferred to the pool, plus the associated fees, by $\mathsf{Flash}_{callback}$. Should the arbitrage prove non-profitable, or the contract fail to pay back the flash loan for any reason, the entire transaction is concluded, resulting in the pool never having transferred the loan in the first place. This is possible due to the entire flash process occurring in a single Ethereum transaction.
@@
Uniswap V3 introduced an NFT-based approach, using an ERC721 wrapper, to track ownership of liquidity positions. This approach allows for a streamlined process for the verification and transfer of ownership of a position. This can be also adopted in $\mathsf{ammBoost}$. At a high level, $\mathsf{TokenBank}$ can be extended to support the NFT approach by utilizing the same implementation found in Uniswap. The caveat though is that creating an NFT will wait until the end of the epoch since it requires mainchain operation (now in $\mathsf{ammBoost}$ positions can be created immediately and synced back to the mainchain since tracking ownership relies on the LPs’ public keys and identifiers are generated at random). Thus, any operations on these new positions has to wait until the next epoch after creating the position NFT.
@@ Appendix 0.C Uniswap Traffic Analysis
In order to find the volume of each transaction type, we used the following query on Dune analytics uniswap_v3_ethereum dataset (the following citation is a direct link to the query used [8]). The query retrieves and counts all of the transactions happened since 2019, splitting them by year and transaction type. uniswap_v3_ethereum is one of Dune analytics ”decoded projects”, meaning that it is a dataset formed from the ABI of the smart contracts operate the protocol in question. Once a user submits a contract for decoding, Dune uses the ABI to generate a table of transactions, query-able by function call or event. As such the uniswap_v3_ethereum is a set of tables which contain the decoded smart contracts which make up the Uniswap V3 protocol. Since we are interested in the transaction volumes, the above query counts all the transactions since 2019 by pulling any row in the tables uniswap_v3_ethereum.Pair_call_burn, uniswap_v3_ethereum.Pair_call_collect, etc. (for swaps, mints, and flashes) for which the call_block_time is $\geq$ 01/01/2019.
Traffic distribution or transaction type frequency. We calculate the frequency of each transaction type by computing the number of transactions for a particular type type divided by the total number of transactions. The volume of transactions is gathered by the query above. The frequency shown in Table 8 is calculated for the year 2023.
@@
Transaction sizes. In order to collect the average size of each transaction type we implemented a python script to interact with an Ethereum node hosted by chainstack[3]. We first collected the transaction hashes of a sufficient amount (approximately 40,000 swaps and 10,000 of all other transaction types) of each transaction type from the Uniswap V3 subgraph [28] (this can similarly be done with the dune query provided above, by modifying the select on line 37 to include the transaction hash). Then we ran a script to analyze the collected transactions. This script basically iterates through the json file which contains our aggregated transaction hashes. For each transaction hash, perform a web3.eth.get_raw_transaction query to retrieve the full raw transaction size. Then, it computes the average size for each transaction type (which is basically sum of total size divided by the number of transactions). The results are also found in Table 8.

### v3

@@ Appendix 0.B Concrete Usecase: Uniswap
Uniswap has three versions: Uniswap V1, released in November 2018, consisted of the baseline protocol that implemented ERC20 token swaps with Ethereum and all of the liquidity management methods (mint, burn, collect). Uniswap V2, released in August 2020, introduced ERC20 to ERC20 swaps, liquidity provision incentives, and oracles. And Uniswap V3 released in May 2021, introduced concentrated liquidity, a nonfungible representation of liquidity positions, and further improvements to the oracle systems. Uniswap adopts the constant product formula for computing the trading price described in Section 2. Uniswap is among the most popular AMMs in practice and commands a large market share in the AMM industry. In this section, we provide an overview of the set of contracts that implement the Uniswap functionality on Ethereum, and the execution trace of the supported transactions.
@@ 0.B.1 Uniswap Supporting Contracts
Pool factory and deployer. The $\mathsf{PoolFactory}$ and $\mathsf{PoolDeployer}$ contracts are responsible for setting up new token pools. The pool deployer contract provides the interface, and the pool factory contract creates the actual pool. Once a pool is created, clients and LPs can start interacting with the pool.
@@ 0.B.2 Transaction Execution Trace
Swap. Regardless of the type of swap (exact in/out) being executed, clients must first deposit their input tokens in the $\mathsf{SwapRouter}$ contract and approve it to spend their tokens. The client then calls the relevant function of the $\mathsf{SwapRouter}$ contract ($\mathsf{ExactInput}$ or $\mathsf{ExactOutput}$) to submit a swap transaction. If the user is performing an $\mathsf{ExactOutput}$ swap, they should implement an additional set of conditional transfers to occur after the call to $\mathsf{ExactOutput}$ to retrieve unspent input tokens.
@@
Burn. To burn a position, the following steps take place:
  • The LP withdraws all the tokens owned by that position. To do so they first call $\mathsf{decreaseLiquidity}$ a method of the
@@
  • The burn function takes the requested amount of tokens to burn, calculates the actual share of liquidity owned by the position which can be burnt (up to the requested amount), and decrements the calculated amount from the positions owned tokens. Finally, it adds the decremented amount to the liquidity positions owed-tokens metric, such that they can be withdrawn by invoking collect.
  • Once the LP have decreased the liquidity owned by their position to zero, they can invoke collect to retrieve those funds before calling $\mathsf{Burn}_{NFPM}$.
  • $\mathsf{Burn}_{NFPM}$ checks that the passed liquidity position does not own any shares of liquidity and all owed tokens have been collected. Should these checks pass it deletes the liquidity position and the NFT associated with it.
Flash. To execute a flash transaction, the client begins by writing and deploying a smart contract which overwrites the $\mathsf{Flash}_{callback}$ method of the liquidity pool. The $\mathsf{Flash}_{callback}$ function is responsible for paying back the loan. As such, should a client want to perform arbitrage with the loan, they begin by overwriting $\mathsf{Flash}_{callback}$, simply adding in solidity code to execute their arbitrage opportunity. They can then call the flash function of the liquidity pool from which they would like to execute a flash transaction. $\mathsf{Flash}$ itself simply transfers the requested loan of tokens to the client, where they are used for the arbitrage opportunity, before being re-transferred to the pool, plus the associated fees, by $\mathsf{Flash}_{callback}$. Should the arbitrage prove non-profitable, or the contract fail to pay back the flash loan for any reason, the entire transaction is concluded, resulting in the pool never having transferred the loan in the first place. This is possible due to the entire flash process occurring in a single Ethereum transaction.
@@
Uniswap V3 introduced an NFT-based approach, using an ERC721 wrapper, to track ownership of liquidity positions. This approach allows for a streamlined process for the verification and transfer of ownership of a position. This can be also adopted in $\mathsf{ammBoost}$. At a high level, $\mathsf{TokenBank}$ can be extended to support the NFT approach by utilizing the same implementation found in Uniswap. The caveat though is that creating an NFT will wait until the end of the epoch since it requires mainchain operation (now in $\mathsf{ammBoost}$ positions can be created immediately and synced back to the mainchain since tracking owenership relies on the LPs’ public keys and identifiers are generated at random). Thus, any operations on these new positions has to wait until the next epoch after creating the position NFT.
@@ Appendix 0.C Uniswap Traffic Analysis
In order to find the volume of each transaction type, we used the following query on Dune analytics uniswap_v3_ethereum dataset (the following citation is a direct link to the query used [10]). The query retrieves and counts all of the transactions happened since 2019, splitting them by year and transaction type. uniswap_v3_ethereum is one of Dune analytics ”decoded projects”, meaning that it is a dataset formed from the ABI of the smart contracts that operate the protocol in question. Once a user submits a contract for decoding, Dune uses the ABI to generate a table of transactions that is query-able by function call or event. As such, the uniswap_v3_ethereum is a set of tables which contain the decoded smart contracts that constitute the Uniswap V3 protocol. Since we are interested in transaction volumes, the above query counts all transactions since 2019 by pulling any row in the tables uniswap_v3_ethereum.Pair_call_burn, uniswap_v3_ethereum.Pair_call_collect, etc., for which the
call_block_time is $\geq$ 01/01/2019.
Traffic distribution or transaction type frequency. We calculate the frequency of each transaction type by computing the number of transactions of that particular type divided by the total number of transactions from all types. The volume of transactions is gathered by the query above. The frequency shown in Table 9 is calculated for the year 2023.
@@
Transaction sizes. In order to collect the average size of each transaction type, we implemented a python script to interact with an Ethereum node hosted by chainstack [4]. We first collected the transaction hashes of a sufficient amount (approximately 40,000 swaps and 10,000 of all other transaction types) of each transaction type from the Uniswap V3 subgraph [29] (this can similarly be done with the dune query provided above, by modifying the select on line 37 to include the transaction hash). Then, we ran a script to analyze the collected transactions. This script basically iterates through the json file which contains our aggregated transaction hashes. For each transaction hash, it performs a web3.eth.get_raw_transaction query to retrieve the full raw transaction size. After that, it computes the average size for each transaction type (which is basically sum of total size divided by the number of transactions). The results are also found in Table 9.